what is the solution of the following system?
-3x-2y=-12
9x + 6y = -9
I believe there is no solution 0=63

Answer by nerdybill(7384)

-3x-2y=-12
9x + 6y = -9
.
multiply top equation by 3 then combine both:
-9x-6y=-36
9x+6y=-9
-----------
0=-45
.
You're correct, in that there is NO solution, however I don't know where you got the 63 from...